# A possible solution to the gallium anomaly moving beyond the leptonic wave function factorization

### v1

@@ Abstract
with $\hat{H}_{\text{GT}}$ denoting the Gamow-Teller Hamiltonian. In order to do so, the approximation that the leptonic wave functions, $\psi_{e,\nu}(\mathbf{r})$ can be factorized out is necessary, invoking the fact that they are expected to be almost constant within the nuclear environment. This allows one to isolate the IBD nuclear matrix element, $\mathcal{M^{\rm IBD}_{\rm nuc}}$, after factorizing the lepton and nuclear parts
@@
Exploiting the detailed-balance principle, it can be equated with that of the inverse reaction, namely
@@
Without invoking the db principle, the ground-state IBD cross section reads as in Eq. (1). Writing out the Gamow-Teller Hamiltonian $\hat{H}_{\rm GT}$, for the specific $^{71}\textrm{Ga}\rightarrow\,^{71}\textrm{Ge}$ process, as well as the leptonic wave functions, we can recalculate the full transition amplitude in Eq. (3) readapting the formalism developed for the $\beta$-decay process in Refs. [63, 64, 65, 66], leading to
@@
This procedure significantly improves the rigor of the neutrino-induced cross-section calculation. However, it requires a precise knowledge of the transition density from nuclear theory calculations. Unfortunately, a first-principles calculation with controlled uncertainties remains challenging [67].
@@
Transition density parametrization.— Here, we assess the impact of the proposed approach using different data-driven phenomenological parametrizations for the transition density. By constraining these models to the precisely measured 71Ge half-life, we demonstrate their potential to resolve the gallium anomaly. In general, each transition density parametrization will depend upon a certain set of parameters $\Theta$, in addition to the radial dependence, such that $\rho_{\rm TD}(r,\Theta)$. We investigate whether (and which) values of $\Theta$ can produce transition densities that lead to a theoretical ground-state cross section which solves the gallium anomaly, i.e. consistent with the values of $\sigma^{\rm exp}_{\text{gs},\,^{51}\mathrm{Cr}}=(4.44\pm 0.33)\times 10^{-45}~\rm{cm}^{2}$ and $\sigma^{\rm exp}_{\text{gs},\,^{37}\mathrm{Ar}}=(5.21\pm 0.72)\times 10^{-45}~\rm{cm}^{2}$, for the 51Cr and 37Ar neutrino sources, respectively. These values have been obtained averaging the experimental results reported by the GALLEX, SAGE, and BEST experiments (as listed in Tab. IV of Ref. [11]) separately for each neutrino source. To isolate the ground-state contribution, we subtract the excited-state component $\Delta_{\text{(p,n)}}$, assigning to it a conservative uncertainty of $\pm 5\%$. [^2] This value is used to include both the $\sim 9\%$ contribution obtained from ($^{3}\text{He},^{3}\text{H}$) reaction data [62] as well as the possible absence of excited state contributions. Based on forward-angle (p,n) scattering data [61], we adopt the values $\Delta_{\text{(p,n)}}=5.3\%$ and 5.8% for the 51Cr and 37Ar sources, respectively.
@@
To find the best set of parameters to solve or alleviate the gallium anomaly, given a certain parametrization, we perform a fit using the least-squares function
@@
To make the model more realistic, we constrain the Gamow-Teller transition density in Eq. (6) so that it reproduces the measured 71Ge electron-capture half-life, using the same nuclear transition density for the inverse 71Ga($\nu_{e}$, $e^{-}$)71Ge reaction (up to a trivial overall phase). The former is related to the electron-capture rate as [59, 68, 69, 48, 70]
where $E^{1s}_{\nu}$ is the neutrino energy corresponding to a $1$s shell electron capture and $P_{\text{L,M,K}}$ are the experimentally measured electron-capture probabilities for the L, M and K shells [52, 71, 59]. [^3] Here, as well as in the IBD cross section, we account for radiative corrections and the weak magnetism contribution following Ref. [37], despite not explicitly written. Here, $\mathcal{H}^{\rm EC}_{1s}$ is the EC matrix element, which for an allowed electron capture involving $1$s shell electrons is
$+\tfrac{1}{3}f_{1\mathrm{s}}(r)j_{1}(q\,r)\big]\Big|^{2}\,$, (12)
with $g_{1\mathrm{s}}$ and $f_{1\mathrm{s}}$ being the large and small bound electron wave-function radial components. In this case, we define the following $\chi^{2}$
@@
Performing a combined fit, obtained by summing the $\chi^{2}$ functions in Eqs. (10) and (13), one may find a transition density parametrization that simultaneously reproduces the experimental half-life of germanium and resolves or alleviates the gallium anomaly by reducing the ground-state cross section up to $\sim 20\%$. To do so, we tested different parametrizations, namely Single Gaussian (SG), Sum of Gaussians (SOG), and Double Gaussian (DG) defined as
$\rho_{\mathrm{TD}}^{\rm SG}(r,\Theta^{\rm SG})=A\,e^{-(r-r_{a})^{2}/2a^{2}}\,$,
where the parameters that are fitted for are $\Theta^{\rm SG}\!=\!\{A,r_{a},a\}$, $\Theta^{\rm SOG}\!=\!\{A_{1},r_{1},A_{2},r_{2},a\}$, and $\Theta^{\rm DG}\!=\!\{A,B,a,b,r_{a},r_{b}\}$, respectively. Moreover, we also tried a variation of the DG shape with the following parameters $\Theta^{\rm DG_{2}}\!=\!\{A,B,a,b,r_{a}=r_{b}\}$. Here, the amplitudes are in units of $\rm fm^{-3}$}], while the radii and standard-deviations in $\rm fm$}]. This choice of parametrizations is guided by ab-initio studies [72, 67] showing that weak transition densities are generally smooth, extending beyond the nuclear radius. However, they may differ in their radial falloff and, for some transitions, display partial cancellations that produce a node or a sign change. In this spirit, SG provides a minimal single-lobe description, DG is the simplest form that can accommodate a sign-changing profile, and SOG offers a straightforward way to add limited shape flexibility when needed without introducing a fully model-dependent functional form.
The results of the combined fit are listed in Tab. 1 and the corresponding transition densities are shown in Fig. 1. In the latter, we also report the outcome obtained by adopting a two-parameter Fermi (2pF) shape,
@@
i.e., by effectively identifying the weak transition density with the nuclear electric charge distribution. While this ansatz may look reasonable at first sight, it is conceptually unjustified because a charge density is a ground-state electromagnetic observable, whereas $\rho_{\rm TD}$ is a weak transition quantity. Notably, even in the case of superallowed Fermi transitions, where the Conserved Vector Current hypothesis relates weak and electromagnetic currents, recent theoretical studies have shown that the charged weak density distinctively deviates from the electric charge distribution, for instance by exhibiting a more pronounced surface-peaking due to nuclear structure effects [73, 74, 75]. Consequently, relying on the charge distribution to describe the weak transition density is conceptually unjustified and potentially inaccurate, a conclusion that holds even more strongly for Gamow-Teller transitions where no underlying symmetry relates the two distributions. We therefore include it only as a benchmark to quantify the impact of this common simplifying assumption. We fix $\Theta^{\rm 2pF}=\{A=0.588,\;c=4.56\,{\rm fm},\;a=0.52\,{\rm fm}\}$: $c$ and $a$ reproduce the experimental 71Ge charge radius, $R_{\rm ch}=4.032\pm 0.002$ fm [58], while $A$ is chosen to match $\mathcal{M}_{\rm nuc}^{\rm EC}$ within the detailed-balance-based approximation. This choice corresponds to the transition-density assumption underlying the latest IBD cross-section prediction in Ref. [11]; hence, that prediction is recovered only insofar as $\rho_{\rm TD}(r)$ is taken to follow the given 2pF shape. In contrast, when confronted with our combined constraints, the resulting ground-state IBD cross section is significantly larger than the experimental one.

### v2

@@ Abstract
with $\hat{H}_{\text{GT}}$ denoting the Gamow-Teller Hamiltonian. In order to do so, the approximation that the leptonic wave functions, $\psi_{e,\nu}(\mathbf{r})$, can be factorized out is necessary, invoking the fact that they are expected to be almost constant within the nuclear environment. This allows one to isolate the IBD nuclear matrix element, $\mathcal{M^{\rm IBD}_{\rm nuc}}$, after factorizing the lepton and nuclear parts
@@
Exploiting the detailed-balance principle, the IBD nuclear matrix element can be equated with that of the inverse reaction, namely
@@
Without invoking the db principle, the ground-state IBD cross section reads as in Eq. (1). Writing out the Gamow-Teller Hamiltonian $\hat{H}_{\rm GT}$, for the specific $^{71}\textrm{Ga}\rightarrow\,^{71}\textrm{Ge}$ process, as well as the leptonic wave functions, we can recalculate the full transition amplitude in Eq. (3) readapting the formalism developed for the $\beta$-decay process in Refs. [63, 64, 65, 66]. Retaining only the dominant axial-spatial component relevant for an allowed transition, one gets
@@
This procedure significantly improves the rigor of the neutrino-induced cross-section calculation. However, it requires a precise knowledge of the transition density from nuclear theory. Unfortunately, a first-principles calculation with controlled uncertainties remains challenging [67].
The same formalism can also be applied to the electron-capture rate, which is related to the germanium half-life as [59, 68, 69, 48, 70]
where $E^{1s}_{\nu}$ is the neutrino energy corresponding to a $1$s shell electron capture and $P_{\text{L,M,K}}$ are the experimentally measured electron-capture probabilities for the L, M and K shells [52, 71, 59]. [^2] Here, as well as in the IBD cross section, we account for radiative corrections and the weak magnetism contribution following Ref. [37], despite not explicitly written. Here, $\mathcal{H}^{\rm EC}_{1s}$ is the EC matrix element, which for an allowed electron capture involving $1$s shell electrons, restricted to the leading axial-spatial component, is
$+\tfrac{1}{3}f_{1\mathrm{s}}(r)j_{1}(q\,r)\big]\Big|^{2}\,$, (11)
with $g_{1\mathrm{s}}$ and $f_{1\mathrm{s}}$ being the large and small bound electron wave-function radial components. Being the EC transition density the same as that of the IBD 71Ga($\nu_{e}$, $e^{-}$)71Ge reaction (up to a trivial overall phase), the EC process serves as a crucial validity test. Indeed, since the 71Ge half-life is experimentally well-constrained, any phenomenological $\rho_{\rm TD}$ capable of resolving the gallium anomaly must simultaneously reproduce the observed germanium half-life.
Here, we assess the impact of the proposed approach using different data-driven phenomenological parametrizations for the transition density. By constraining these models to the precisely measured 71Ge half-life, we demonstrate their potential to resolve the gallium anomaly.
@@
Transition density parametrization.— In general, each transition density parametrization will depend upon a certain set of parameters $\Theta$, in addition to the radial dependence, such that $\rho_{\rm TD}(r,\Theta)$. We investigate whether (and which) values of $\Theta$ can produce transition densities that lead to a theoretical ground-state cross section which solves the gallium anomaly, i.e. consistent with the values of $\sigma^{\rm exp}_{\text{gs},\,^{51}\mathrm{Cr}}=(4.44\pm 0.33)\times 10^{-45}~\rm{cm}^{2}$ and $\sigma^{\rm exp}_{\text{gs},\,^{37}\mathrm{Ar}}=(5.21\pm 0.72)\times 10^{-45}~\rm{cm}^{2}$, for the 51Cr and 37Ar neutrino sources, respectively. These values have been obtained averaging the experimental results reported by the GALLEX, SAGE, and BEST experiments (as listed in Tab. IV of Ref. [11]) separately for each neutrino source. To isolate the ground-state contribution, we subtract the excited-state component $\Delta_{\text{(p,n)}}$, assigning to it a conservative uncertainty of $\pm 5\%$. [^3] This value is used to include both the $\sim 9\%$ contribution obtained from ($^{3}\text{He},^{3}\text{H}$) reaction data [62] as well as the possible absence of excited state contributions. Based on forward-angle (p,n) scattering data [61], we adopt the values $\Delta_{\text{(p,n)}}=5.3\%$ and 5.8% for the 51Cr and 37Ar sources, respectively.
In this work, we tested different parametrizations, namely Single Gaussian (SG), Sum of Gaussians (SOG), and Double Gaussian (DG) defined as
$\rho_{\mathrm{TD}}^{\rm SG}(r,\Theta^{\rm SG})=A\,e^{-(r-r_{a})^{2}/2a^{2}}\,$,
where the parameters that are fitted for are $\Theta^{\rm SG}\!=\!\{A,r_{a},a\}$, $\Theta^{\rm SOG}\!=\!\{A_{1},r_{1},A_{2},r_{2},a\}$, and $\Theta^{\rm DG}\!=\!\{A,B,a,b,r_{a},r_{b}\}$, respectively. Moreover, we also tried a variation of the DG shape, referred to as $\rm DG_{2}$}, with the following parameters $\Theta^{\rm DG_{2}}\!=\!\{A,B,a,b,r_{a}=r_{b}\}$. Here, the amplitudes are in units of $\rm fm^{-3}$}], while the radii and standard-deviations in $\rm fm$}]. This choice of parametrizations is guided by ab-initio studies [72, 67] showing that weak transition densities are generally smooth, extending beyond the nuclear radius. However, they may differ in their radial falloff and, for some transitions, display partial cancellations that produce a node or a sign change [73]. In this spirit, SG provides a minimal single-lobe description, DG is the simplest form that can accommodate a sign-changing profile, and SOG offers a straightforward way to add limited shape flexibility when needed without introducing a fully model-dependent functional form.
@@
Fit results.— To find the best set of parameters to solve or alleviate the gallium anomaly, given a certain parametrization, we perform a fit using the least-squares function
@@
To make the model more realistic, we constrain the Gamow-Teller transition density in Eq. (6) so that it reproduces the measured 71Ge electron-capture half-life. In this case, we define the following $\chi^{2}$
@@
Performing a combined fit, obtained by summing the $\chi^{2}$ functions in Eqs. (13) and (14), one may find a transition density parametrization that simultaneously reproduces the experimental half-life of germanium and resolves or alleviates the gallium anomaly by reducing the ground-state cross section up to $\sim 20\%$. The results of the combined fit are listed in Tab. 1 and the corresponding transition densities are shown in Fig. 1. In the latter, we also report the outcome obtained by adopting a two-parameter Fermi (2pF) shape [74]
@@
i.e., by effectively identifying the weak transition density with the nuclear electric charge distribution. While this ansatz may look reasonable at first sight, it is conceptually unjustified because a charge density is a ground-state electromagnetic observable, whereas $\rho_{\rm TD}$ is a weak transition quantity. Notably, even in the case of superallowed Fermi transitions, where the Conserved Vector Current hypothesis relates weak and electromagnetic currents, recent theoretical studies have shown that the charged weak density distinctively deviates from the electric charge distribution, for instance by exhibiting a more pronounced surface-peaking due to nuclear structure effects [75, 76, 77]. Consequently, relying on the charge distribution to describe the weak transition density is conceptually unjustified and potentially inaccurate, a conclusion that holds even more strongly for Gamow-Teller transitions where no underlying symmetry relates the two distributions. We therefore include it only as a benchmark to quantify the impact of this common simplifying assumption. We fix $\Theta^{\rm 2pF}=\{A=0.588,\;c=4.56\,{\rm fm},\;a=0.52\,{\rm fm}\}$: $c$ and $a$ reproduce the experimental 71Ge charge radius, $R_{\rm ch}=4.032\pm 0.002$ fm [58], while $A$ is chosen to match $\mathcal{M}_{\rm nuc}^{\rm EC}$ within the detailed-balance-based approximation. This choice corresponds to the transition-density assumption underlying the latest IBD cross-section prediction in Ref. [11]; hence, that prediction is recovered only insofar as $\rho_{\rm TD}(r)$ is taken to follow the given 2pF 71Ge charge density. In contrast, when confronted with our combined constraints, the resulting ground-state IBD cross section is significantly larger than the experimental one.